# Towards Universal Mesh Movement Networks

### v2

@@ Abstract
Solving complex Partial Differential Equations (PDEs) accurately and efficiently is an essential and challenging problem in all scientific and engineering disciplines. Mesh movement methods provide the capability to improve the accuracy of the numerical solution without increasing the overall mesh degree of freedom count. Conventional sophisticated mesh movement methods are extremely expensive and struggle to handle scenarios with complex boundary geometries. However, existing learning-based methods require re-training from scratch given a different PDE type or boundary geometry, which limits their applicability, and also often suffer from robustness issues in the form of inverted elements. In this paper, we introduce the Universal Mesh Movement Network (UM2N), which – once trained – can be applied in a non-intrusive, zero-shot manner to move meshes with different size distributions and structures, for solvers applicable to different PDE types and boundary geometries. UM2N consists of a Graph Transformer (GT) encoder for extracting features and a Graph Attention Network (GAT) based decoder for moving the mesh. We evaluate our method on advection and Navier-Stokes based examples, as well as a real-world tsunami simulation case. Our method outperforms existing learning-based mesh movement methods in terms of the benchmarks described above. In comparison to the conventional sophisticated Monge-Ampère PDE-solver based method, our approach not only significantly accelerates mesh movement, but also proves effective in scenarios where the conventional method fails. Our project page is at https://erizmr.github.io/UM2N/.
@@ 1 Introduction
An alternative approach to improve the efficiency of a PDE solver is to utilize mesh adaptation, which is a technique for distributing mesh according to the requirements of numerical accuracy. Two main categories of mesh adaptation techniques can be identified: $h$-adaptation and $r$-adaptation. $h$-adaptation refines or coarsens the mesh resolution dynamically through topological operations such as adding/deleting nodes and swapping element edges/faces. In contrast, $r$-adaptation (or mesh movement) relocates or moves mesh nodes, keeping the mesh topology and thus the number of elements and vertices in the mesh unchanged [9]. These traditional mesh adaptation techniques can help reduce PDE-solving costs, but the mesh adaptation process itself may come at the cost of significant computational overhead.
Deep learning based methods have been proposed to accelerate mesh adaptation. [10, 11, 12, 13, 14, 15]. Most previous methods focus on $h$-adaptive refinement [11, 12, 13, 16]. Some existing works focus on learning error indicators or Riemannian metrics (which account for element shape and anisotropy, as well as size) [13, 14, 15] to guide the mesh adaptation process. The learned indicator or metric is then fed into a traditional remesher for mesh generation, or a mesh optimizer, which overall limits the performance of these works to being no better than traditional methods, especially in terms of efficiency. Reinforcement learning based methods show potential to improve the mesh adaptation task, but are difficult to train with low data efficiency [11, 16, 17]. Only a small number of works focus on $r$-adaptation. [18] and [19] investigated supervised and unsupervised learning based mesh movement methods. However, the proposed methods require re-training from scratch given a different PDE or geometry, which limits their applicability. In addition, the unsupervised learning method requires significant GPU memory and suffers from long training times, which makes re-training even more prohibitive for large scale problems.
@@
$\bullet$ We present the UM2N framework, which once trained can be applied in a non-intrusive, zero-shot manner to move meshes with different sizes and structures, for solvers applicable to different PDE types and boundary geometries.
$\bullet$ We propose PDE-independent training strategies and mesh tangling aware loss functions for our Graph Transformer and GAT-based architecture.
$\bullet$ We demonstrate the universal ability of UM2N on different PDEs and boundary geometries: Advection, Navier-Stokes examples, and a real-world tsunami simulation case.
@@ 4 UM2N: Universal Mesh Movement Network
Graph Attention Network (GAT) based decoder. We choose GATs to construct our decoder. Its attention mechanism can help constrain the mesh vertex movement to within one hop of its neighbors which assists in alleviation of mesh tangling issues. The decoder consists of $N$ GAT blocks. The first GAT block takes a mesh query $\boldsymbol{\xi}$, and the embedding $\boldsymbol{z}$ from the transformer encoder as well as the edge features $\boldsymbol{e}$ as inputs. For each following $k$th block, the inputs consist of the coordinates of the initial mesh $\boldsymbol{\xi}$, the coordinates of the intermediate moved mesh $\boldsymbol{\xi}^{(k-1)}$ and extracted features from the previous layer $\boldsymbol{f}^{(k-1)}$.
@@
We denote the modified (adapted) mesh with model parameters $\boldsymbol{\theta}$ as $\mathcal{M}(\boldsymbol{\xi},m_{\boldsymbol{\xi}};\boldsymbol{\theta})=\{\boldsymbol{V}_{a},\boldsymbol{T}_{a}\}$, where $\boldsymbol{V}_{a}=\{\boldsymbol{x}_{i}|i=1,2,\ldots,n_{1}\}$ represents the coordinates of vertices and $\boldsymbol{T}_{a}=\{\boldsymbol{t}_{i}|i=1,2,\ldots,n_{2}\}$ represents the elements with $\boldsymbol{x}_{i},\boldsymbol{t}_{i}$ as the $i$-th vertex and element. For ease of presentation, we omit the $\boldsymbol{\theta}$ here. Similarly, we define $\boldsymbol{y}_{i},\boldsymbol{q}_{i}$ as $\boldsymbol{V}_{r},\boldsymbol{T}_{r}$’s $i$-th vertex and element.
@@
$L_{vol}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{\xi},m_{\boldsymbol{\xi}};\boldsymbol{V}_{r},\boldsymbol{T}_{r})\in\mathcal{D}}\left[\frac{1}{|\boldsymbol{T}_{a}|}\sum_{i=1}^{n_{2}}|\text{Vol}(\boldsymbol{t}_{i})-\text{Vol}(\boldsymbol{q}_{i})|\right],$ (6)
@@
$L_{cd}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{\xi},m_{\boldsymbol{\xi}};\boldsymbol{V}_{r},\boldsymbol{T}_{r})\in\mathcal{D}}\left[\frac{1}{|\boldsymbol{V}_{a}|}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{V}_{a}}\min_{\boldsymbol{y}_{j}\in\boldsymbol{V}_{r}}\|\boldsymbol{x}_{i}-\boldsymbol{y}_{j}\|_{\ell_{2}}+\frac{1}{|\boldsymbol{V}_{r}|}\sum_{y_{j}\in\boldsymbol{V}_{r}}\min_{\boldsymbol{x}_{i}\in\boldsymbol{V}_{a}}\|\boldsymbol{x}_{i}-\boldsymbol{y}_{j}\|_{\ell_{2}}\right].$ (7)
@@ 5 Experiments
Unlike existing approaches to mesh movement, such as [18, 19], which re-train their model on a case-by-case basis, our study aims to explore the universal applicability of the proposed UM2N. Therefore, all examples shown in this section are tested in a zero-shot generalization manner, i.e., without re-training. In addition to PDE type, our training meshes are of small size typically comprising $500$ vertices, while our test meshes have far more vertices: Advection ($2,052$ vertices), Cylinder ($4,993$ vertices), Tsunami ($8,117$ vertices), which also tests the model’s scalability. There are examples with more complicated settings, please refer to Appx. D.
@@
In the following section, we benchmark the mesh movement using both non-learned Monge-Ampère (MA) mesh movement implemented in Movement [39] and learning-based baseline M2N [18] over three different scenarios to compare with UM2N. Table 1 shows the quantitative results in three different scenarios, Swirl, Cylinder, and Helmholz. An additional qualitative result of Tōhoku Tsunami simulation, is provided to show our model can handle highly complex boundaries in real-world scenarios. Moreover, to give a fair comparison, we re-trained the model in M2N on our training dataset before evaluation.
Helmholtz. The Helmholtz case is a time-independent problem solving a 2D Helmholtz equation as shown in (A2). The source terms of the equation are generated use formula as shown in (A1). Our model achieves better error reduction than M2N and comparable to MA method with significant less time cost.
@@
The results are shown in Fig. 2. As shown in Table 1, our UM2N shows a average $35.93\%$ error reduction which clearly outperforms M2N’s $14.15\%$. It is close to that of the full Monge-Ampère (MA) mesh movement approach which achieves an error reduction of $37.43\%$, but with significantly lower computational time.
@@
We investigate the dynamics in the area around the cylinder and in the wake flow respectively. For the former, we use the measure of $C_{D}$ as shown in Fig. 3. The results for $C_{L}$ is shown in Fig. A5 in Appx. D. For the wake flow, we use cross-section probes at $x=[0.5,1.0,1.5,2.0]$ to measure the intensity of vorticity spatially along the $y$-axis as shown in Fig. 4.
@@
Note that we intended to compare our UM2N to M2N and the conventional full MA method. However, both MA and M2N fail i.e., the solver diverges, within 30 time steps due to mesh tangling and thus provide insufficient data points for comparison.
@@
5.3 Ablation study.
Volume loss vs coordinate loss. We denote the UM2N variant trained with coordinate loss [18] as UM2N-coord. As shown in Table 2, UM2N-coord achieves comparable error reduction ($31.21\%$) to the proposed UM2N ($35.93\%$) on the Swirl case. However, it fails (i.e., encounters mesh tangling) on the Cylinder case which has a more complex boundary geometry. This indicates that the element volume loss is mesh-tangling aware, i.e., helps to prevent producing inverted elements.
@@ 6 Conclusions
We introduce the Universal Mesh Movement Network (UM2N), which once trained on our generated PDE-independent dataset, can be applied in a non-intrusive, zero-shot manner to move meshes with different sizes and structure, for solvers applicable to different PDE types and boundary geometries. Our method demonstrates superior mesh movement results on Advection and Navier-Stokes PDE examples as well as a real-world tsunami simulation case. Our UM2N outperforms existing learning based method and achieves comparable PDE error reduction to the costly Monge-Ampére (MA) PDE based approach. Our UM2N demonstrates effectiveness on cases with complex boundary geometries where existing learning and conventional MA based methods fail. Our method also shows significant acceleration compared to MA based mesh adaptation. See Appx. F for the discussions on limitations and broader impacts.
@@
All training data are generated within a 2-D unit square domain. Equivalent sizes $n\times n$ meshes ($n\in\{18,20\}$) are generated for model training. Here we show samples of the generated fields in Fig. A1.
@@
Appendix C Details of experimental settings
C.1 Helmholtz
@@ C.2 Swirl
Monitor Function
$m({\bf x})=1+\max\left(\alpha\frac{\|H(u)\|}{\max\|H(u)\|},\beta\frac{\|G(u)\|}{\max\|G(u)\|}\right)$ (A5)
The $\|H(u)\|$ and $\|G(u)\|$ are hessian norm and gradient norm respectively. In practical, we set $\alpha=5$ and $\beta=10$.
C.3 Flow past a cylinder
@@
Appendix D Additional experimental results.
@@
Appendix E Mesh movement enhanced simulation.
@@
Appendix F Limitations and broader impacts.
Limitations. 1) The GAT to an extent constrains the mesh movement to lie in a restricted range, which may be less effective in scenarios requiring substantial large deformation within a single iteration of the mesh mover. 2) There is no theoretical guarantee that mesh tangling will be completely prevented. 3) Given a low-quality or extremely coarse original mesh, the efficacy of the proposed method may be constrained by the inherent limitation on the degree of freedom count, as it reallocates rather than increases the number of mesh vertices. A promising future direction could be exploring the integration of mesh movement with $h$-adaptation, potentially developing a hybrid $hr$-adaptation approach that combines the advantages of both techniques.

### v3

@@ Abstract
Solving complex Partial Differential Equations (PDEs) accurately and efficiently is an essential and challenging problem in all scientific and engineering disciplines. Mesh movement methods provide the capability to improve the accuracy of the numerical solution without increasing the overall mesh degree of freedom count. Conventional sophisticated mesh movement methods are extremely expensive and struggle to handle scenarios with complex boundary geometries. However, existing learning-based methods require re-training from scratch given a different PDE type or boundary geometry, which limits their applicability, and also often suffer from robustness issues in the form of inverted elements. In this paper, we introduce the Universal Mesh Movement Network (UM2N), which – once trained – can be applied in a non-intrusive, zero-shot manner to move meshes with different size distributions and structures, for solvers applicable to different PDE types and boundary geometries. UM2N consists of a Graph Transformer (GT) encoder for extracting features and a Graph Attention Network (GAT) based decoder for moving the mesh. We evaluate our method on advection and Navier-Stokes based examples, as well as a real-world tsunami simulation case. Our method out-performs existing learning-based mesh movement methods in terms of the benchmarks described above. In comparison to the conventional sophisticated Monge-Ampère PDE-solver based method, our approach not only significantly accelerates mesh movement, but also proves effective in scenarios where the conventional method fails. Our project page can be found at https://erizmr.github.io/UM2N/.
@@ 1 Introduction
An alternative approach to improve the efficiency of a PDE solver is to utilize mesh adaptation, which is a technique for distributing the mesh according to numerical accuracy requirements. Two main categories of mesh adaptation techniques can be identified: $h$-adaptation and $r$-adaptation. $h$-adaptation refines or coarsens the mesh resolution dynamically through topological operations such as adding/deleting nodes and swapping element edges/faces. In contrast, $r$-adaptation (or mesh movement) relocates or moves mesh nodes, keeping the mesh topology and thus the number of elements and vertices in the mesh unchanged [9]. These traditional mesh adaptation techniques can help reduce PDE-solving costs, but the mesh adaptation process itself may come at the cost of significant computational overhead.
Deep learning based methods have been proposed to accelerate mesh adaptation. [10, 11, 12, 13, 14, 15, 16]. Most previous methods focus on $h$-adaptive refinement [11, 12, 13, 17]. Some existing works focus on learning error indicators or Riemannian metrics (which account for element shape and anisotropy, as well as size) [13, 14, 15, 16] to guide the mesh adaptation process. The learned indicator or metric is then fed into a traditional remesher for mesh generation, or a mesh optimizer, which overall limits the performance of these works to being no better than traditional methods, especially in terms of efficiency. Reinforcement learning based methods show potential to improve the mesh adaptation task, but are difficult to train with low data efficiency [11, 17, 18]. Only a small number of works focus on $r$-adaptation. [19, 20] investigated supervised and unsupervised learning based mesh movement methods. However, the proposed methods require re-training from scratch given a different PDE or geometry, which limits their applicability. In addition, the unsupervised learning method requires significant GPU memory and suffers from long training times, which makes re-training even more prohibitive for large scale problems.
@@
  • We present the UM2N framework, which once trained can be applied in a non-intrusive, zero-shot manner to move meshes with different sizes and structures, for solvers applicable to different PDE types and boundary geometries.
  • We propose PDE-independent training strategies and mesh tangling aware loss functions for our Graph Transformer and GAT-based architecture.
  • We demonstrate the universal ability of UM2N on different PDEs and boundary geometries: Advection, Navier-Stokes examples, and a real-world tsunami simulation case.
@@ 4 UM2N: Universal Mesh Movement Network
Graph Attention Network (GAT) based decoder. We choose GATs to construct our decoder. Its attention mechanism can help constrain the mesh vertex movement to within one hop of its neighbors which assists in alleviation of mesh tangling issues. The decoder consists of $N$ GAT blocks. The first GAT block takes a mesh query $\boldsymbol{\xi}$, and the embedding $\boldsymbol{z}$ from the transformer encoder as well as the edge features $\boldsymbol{e}$ as inputs. For each following $k^{\text{th}}$ block, the inputs consist of the coordinates of the initial mesh $\boldsymbol{\xi}$, the coordinates of the intermediate moved mesh $\boldsymbol{\xi}^{(k-1)}$ and extracted features from the previous layer $\boldsymbol{f}^{(k-1)}$.
@@
We denote the modified (adapted) mesh with model parameters $\boldsymbol{\theta}$ as $\mathcal{M}(\boldsymbol{\xi},m_{\boldsymbol{\xi}};\boldsymbol{\theta})=\{\boldsymbol{V}_{a},\boldsymbol{T}_{a}\}$, where $\boldsymbol{V}_{a}=\{\boldsymbol{x}_{i}\}_{i=1}^{n_{1}}$ represents the coordinates of vertices and $\boldsymbol{T}_{a}=\{\boldsymbol{t}_{i}\}_{i=1}^{n_{2}}$ represents the elements with $\boldsymbol{x}_{i},\boldsymbol{t}_{i}$ as the $i^{\text{th}}$ vertex and element. For ease of presentation, we omit the $\boldsymbol{\theta}$ here. Similarly, we define $\boldsymbol{y}_{i},\boldsymbol{q}_{i}$ as $\boldsymbol{V}_{r},\boldsymbol{T}_{r}$’s $i^{\text{th}}$ vertex and element.
@@
$L_{vol}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{\xi},m_{\boldsymbol{\xi}};\boldsymbol{V}_{r},\boldsymbol{T}_{r})\in\mathcal{D}}\left[\frac{1}{|\boldsymbol{T}_{a}|}\sum_{i=1}^{n_{2}}\left|\text{Vol}(\boldsymbol{t}_{i})-\text{Vol}(\boldsymbol{q}_{i})\right|\right],$ (6)
@@
$L_{cd}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{\xi},m_{\boldsymbol{\xi}};\boldsymbol{V}_{r},\boldsymbol{T}_{r})\in\mathcal{D}}\left[\frac{1}{|\boldsymbol{V}_{a}|}\sum_{\boldsymbol{x}_{i}\in\boldsymbol{V}_{a}}\min_{\boldsymbol{y}_{j}\in\boldsymbol{V}_{r}}\|\boldsymbol{x}_{i}-\boldsymbol{y}_{j}\|_{2}+\frac{1}{|\boldsymbol{V}_{r}|}\sum_{y_{j}\in\boldsymbol{V}_{r}}\min_{\boldsymbol{x}_{i}\in\boldsymbol{V}_{a}}\|\boldsymbol{x}_{i}-\boldsymbol{y}_{j}\|_{2}\right].$ (7)
@@ 5 Experiments
Unlike existing approaches to mesh movement, such as [19, 20], which re-train their model on a case-by-case basis, our study aims to explore the universal applicability of the proposed UM2N. Therefore, all examples shown in this section are tested in a zero-shot generalization manner, i.e., without re-training. In addition to PDE type, our training meshes are of small size typically comprising $500$ vertices, while our test meshes have far more vertices: Advection ($2,052$ vertices), Cylinder ($4,993$ vertices), Tsunami ($8,117$ vertices), which also tests the model’s scalability. There are examples with more complicated settings, please refer to Appx. E.
@@
In the following section, we benchmark the mesh movement using both non-learned Monge-Ampère (MA) mesh movement implemented in Movement [40] and learning-based baseline M2N [19] over three different scenarios to compare with UM2N. Table 1 shows the quantitative results in three different scenarios, Swirl, Cylinder, and Helmholtz. An additional qualitative result of Tōhoku Tsunami simulation, is provided to show our model can handle highly complex boundaries in real-world scenarios. Moreover, to give a fair comparison, we re-trained the model in M2N on our training dataset before evaluation.
Helmholtz. The Helmholtz case is a time-independent problem solving a 2D Helmholtz equation as in (A2). The source terms of the equation are generated use formula (A1). Our model achieves better error reduction than M2N and comparable to MA method in significantly less time.
@@
The results are shown in Fig. 2. As shown in Table 1, our UM2N shows an average $35.93\%$ error reduction which clearly outperforms M2N’s $14.15\%$. It is close to that of the full Monge-Ampère (MA) mesh movement approach which achieves an error reduction of $37.43\%$, but with significantly lower computational time.
@@
We investigate the dynamics in the area around the cylinder and in the wake flow respectively. For the former, we use the measure of $C_{D}$ as shown in Fig. 3. The results for $C_{L}$ is shown in Fig. A10 in Appx. E. For the wake flow, we use cross-section probes at $x=[0.5,1.0,1.5,2.0]$ to measure the intensity of vorticity spatially along the $y$-axis as shown in Fig. 4.
@@
Note that we intended to compare our UM2N to M2N and the conventional full MA method. However, both MA and M2N fail i.e., the solver diverges, within 30 time steps due to mesh tangling and thus provide insufficient data points for comparison. For comparisons with a non-tangled M2N case, please refer to the simplified laminar flow-past-cylinder example presented in Appendix D.1.
@@
Figure 7: Tangled Case of UM2N and MA. Red/blue indicate high/low monitor values, i.e., the mesh is expected to stretch towards the right direction. The case is more and more challenging from top to bottom.
5.3 Tangled results of the UM2N model
There are in fact still cases where UM2N might fail. To investigate these, we perform additional experiments as shown in Figure 7. Starting from a rectangle, we gradually distort the geometry to a more and more non-convex shape, i.e., the case becomes more challenging from top to bottom. The UM2N finally fails in the extreme case at sudden jumps in the boundary accompanied by a large variation in the required resolution as shown in the bottom row of Figure R1: a sudden constriction of the flow leads to tangling with UM2N in the two left corners of the channel. In less extreme cases, UM2N does produce a valid mesh, whereas the original MA method still fails (as seen in the second to bottom rows in Figure 7).
5.4 Ablation study
@@
Volume loss vs coordinate loss.
We denote the UM2N variant trained with coordinate loss [19] as UM2N-coord. As shown in Table 2, UM2N-coord achieves comparable error reduction ($31.21\%$) to the proposed UM2N ($35.93\%$) on the Swirl case. However, it fails (i.e., encounters mesh tangling) on the Cylinder case which has a more complex boundary geometry. This indicates that the element volume loss is mesh-tangling aware, i.e., helps to prevent producing inverted elements.
@@ 6 Conclusions
We introduce the Universal Mesh Movement Network (UM2N), which once trained on our generated PDE-independent dataset, can be applied in a non-intrusive, zero-shot manner to move meshes with different sizes and structures, for solvers applicable to different PDE types and boundary geometries. Our method demonstrates superior mesh movement results on Advection and Navier-Stokes PDE examples as well as a real-world tsunami simulation case. Our UM2N outperforms the existing learning based method and achieves comparable PDE error reduction to the far more costly Monge-Ampére (MA) PDE based approach. Our UM2N demonstrates effectiveness in cases with complex boundary geometries where existing learning and conventional MA based methods fail. Our method also shows significant acceleration compared to MA based mesh adaptation. See Appx. G for discussions on limitations and broader impacts.
Acknowledgment
Thanks to Jinghan Jia for testing UM2N performance on Nvidia A100.
@@ Acknowledgment
We include additional experimental details and results in the appendix. Detailed information on the software used is provided in Appendix A. A description of the dataset generation process is presented in Appendix B. We provide the details of experimental settings of cases in the paper in Appendix C. To provide a more holistic analysis for the proposed UM2N, we include evaluations across various test cases, performance benchmarks, more ablation studies, and assessments on low-quality initial meshes in Appendix D. To demonstrate the the universal ability of the UM2N, additional experiments conducted under diverse settings are detailed in Appendix E. We introduce the mesh movement-enhanced simulation, illustrating the non-intrusive integration of UM2N into a numerical simulation pipeline in Appendix F. A discussion about the limitations and broader impacts of our approach is provided in Appendix G.
@@
The code used to produce the results presented in this paper is publicly available at https://github.com/mesh-adaptation/UM2N. (Release [49].) Code documentation is hosted at https://mesh-adaptation.github.io.
@@
All training data are generated within a 2-D unit square domain. Equivalent size $n\times n$ meshes ($n\in\{18,20\}$) are generated for model training. Here we show samples of the generated fields in Fig. A1.
@@
Appendix C Details of experimental settings.
C.1 Helmholtz.
@@
Monitor Function.
$m({\bf x})=1+\max\left(\alpha\frac{\|H(u)\|}{\max\|H(u)\|},\beta\frac{\|G(u)\|}{\max\|G(u)\|}\right),$ (A5)
where we set $\alpha=5$ and $\beta=10$. The $\|H(u)\|$ and $\|G(u)\|$ are Hessian norm and gradient norm respectively.
C.3 Flow past a cylinder.
@@
Appendix D More Evaluations.
Figure A2: Non-tangled results of M2N on flow-past-a-cylinder case.
Table A1: Measure between UM2N and M2N model output mesh on the flow-past-a-cylinder in Fig. A2.
UM2N M2N
Property ER $\uparrow$ (%) ER $\uparrow$ (%)
Vorticity 27.43 % 10.22 %
D.1 Non-tangled results of other methods on the flow-past-cylinder.
We additionally perform a non-tangled flow-past-cylinder shown in Fig. A2. We simplified the setting: lower the Rrenynolds number by reducing the inflow velocity and set the top and bottom slip boundary, which finally makes it a laminar flow. As shown in Fig A2, both the UM2N and M2N give non-tangled meshes and it can be observed that UM2N better captures the dynamics i.e., adapts to the PDE solution of stable state compared to the High Res solution. The quantitative results shown in the Table A1 also indicate that UM2N perform better than M2N regarding to error reduction.
D.2 Performance benchmark.
To evaluate the runtime performance of UM2N, we constructed a benchmark case based on the Swirl scenario, consisting of 11,833 vertices and 23,668 triangular elements. We extended the one-ring setup to a nine-ring configuration. In this benchmark, we compare the UM2N output mesh and the reference mesh generated by the MA method, as well as the solutions obtained on these meshes, against those computed on high-resolution and original meshes as shown in Fig. A3.
We tested the inference time of UM2N on two GPUs (Nvidia GeForce RTX 3090 and Nvidia A100) and compared its performance to the MA method. For a fair comparison, we also included CPU benchmarks, as the current state-of-the-art MA implementation is limited to CPU execution. The CPU used for testing was an 11th Gen Intel(R) Core(TM) i9-11900K @ 3.50GHz. The performance results are presented in Table A2.
The results indicate that UM2N achieved more than 2000x speedup on the Nvidia RTX 3090 and over 3000x on the Nvidia A100, compared to the MA method. Even on the CPU, UM2N demonstrated an approximate 100x speedup. In summary, UM2N provides comparable solution accuracy while delivering a significant performance improvement.
Figure A3: Visualization of UM2N output mesh and reference mesh from MA. Comparison between solutions obtained on high resolution, original, UM2N and MA meshes.
Table A2: Performance benchmark on test case shown in Figure. A3.
MA UM2N UM2N UM2N
Hardware CPU CPU GPU (Nvida RTX 3090) GPU (Nvida A100)
Avg. inference time (s) $114.3$ $1.113$ $0.0487$ $0.0342$
Speedup - 102$\sim$ 102$\times$ 102 × 2347$\sim$ 2347$\times$ 2347 × 3342$\sim$ 3342$\times$ 3342 ×
D.3 Boundary layer case.
Figure A4: Boundary Layer. The two plots at left visualize the $u_{y}$ from boundary to $y=0.1$. It can be observed that UM2N improves the accuracy of the velocity obtained on uniform mesh compared to High Res solution.
In our flow past the cylinder case, at the top and bottom we imposed non-slip boundary conditions which already promotes a boundary layer similar to the suggested flow over a flat plate. For a clearer illustration, we have now additionally conducted a flow past two parallel plates experiment. We observe that the velocity parallel to the boundary ($u_{y}$) obtained from UM2N moved mesh better aligns with the high resolution results compared to the results on uniform mesh as shown in the Fig. A4.
D.4 Poor quality initial mesh.
Figure A5: A very poor initial mesh case. UM2N perform mesh movement for an anisotropic Helmholtz problem.
We include an additional experiment starting from a poor quality initial mesh as shown in Fig. A5. The input mesh has highly anisotropic elements and several vertices with valence 6. This can certainly be considered a “low quality” mesh. Consider the monitor function to be the $L^{2}$-norm of the recovered gradient of the solution of an anisotropic Helmholtz problem. With this monitor function, we tried applying conventional MA solvers and found that both the quasi-Newton and relaxation approaches failed to converge and/or resulted in tangled meshes, even though the elements are already aligned with the anisotropy of the Helmholtz solution. The UM2N approach, however, was able to successfully apply mesh movement without tangling. We would also like to clarify that our work targeting mesh movement that dynamically adapts to a PDE solution, which is not a replacement for mesh generation in general, so we would always expect a reasonable input mesh when applying our model.
D.5 Further Ablation studies.
We perform an additional ablation study of our network architecture on the Swirl case. We construct two models, which replace the GAT Decoder with one layer GAT denoted as UM2N-w/o-Decoder and remove the graph transformer denoted as UM2N-w/o-GT. The results are shown in the Table A6.It can be observed that both variants give worse results compared to the full UM2N. A visualization is also shown in Fig. A6. It can be observed that without the Decoder, the model distorts the shape of the ring, i.e., missing relative information between vertices; without the graph transformer, the model fails to capture the details of the ring shape.
Swirl
Method ER $\uparrow$ (%)
UM2N-w/o-Decoder 10.68
UM2N-w/o-GT 12.35
UM2N (Ours) 35.93
MA (reference) 37.43
(a) Error reduction of UM2N variants on Swirl case.
(b) Moved from UM2N variants and reference mesh from MA on Swirl case.
Figure A6: Ablation of network components. Full UM2N gives the best error reduction.
Appendix E Additional experimental results.
@@
Appendix F Mesh movement enhanced simulation.
@@
Appendix G Limitations and broader impacts.
Limitations. 1) The GAT to an extent constrains the mesh movement to lie in a restricted range, which may be less effective in scenarios requiring substantial large deformation within a single iteration of the mesh mover. 2) There is no theoretical guarantee that mesh tangling will be completely prevented. 3) The model’s best performance relies on a manually chosen monitor function, though a gradient based monitor function can yield reasonable results. 3) Given a low-quality or extremely coarse original mesh, the efficacy of the proposed method may be constrained by the inherent limitation on the degree of freedom count, as it redistributes rather than increases the number of mesh vertices. A promising future direction could be exploring the integration of mesh movement with $h$-adaptation, potentially developing a hybrid $hr$-adaptation approach that combines the advantages of both techniques.